# ExLLM: Experience-Enhanced LLM Optimization for Molecular Design and Beyond

### v4

@@ Abstract
Molecular design involves an enormous and irregular search space, where traditional optimizers such as Bayesian optimization, genetic algorithms, and generative models struggle to leverage expert knowledge or handle complex feedback. Recently, LLMs have been used as optimizers, achieving promising results on benchmarks such as PMO. However, existing approaches rely only on prompting or extra training, without mechanisms to handle complex feedback or maintain scalable memory. In particular, the common practice of appending or summarizing experiences at every query leads to redundancy, degraded exploration, and ultimately poor final outcomes under large-scale iterative search. We introduce ExLLM (Experience-Enhanced LLM optimization), an LLM-as-optimizer framework with three components: (1) a compact, evolving experience snippet tailored to large discrete spaces that distills non-redundant cues and improves convergence at low cost; (2) a simple yet effective k-offspring scheme that widens exploration per call and reduces orchestration cost; and (3) a lightweight feedback adapter that normalizes objectives for selection while formatting constraints and expert hints for iteration. ExLLM sets new state-of-the-art results on PMO and generalizes strongly—in our setup, it sets records on circle packing and stellarator design, and yields consistent gains across additional domains—requiring only a task-description template and evaluation functions to transfer.
@@ 1 Introduction
We propose ExLLM, an LLM-as-optimizer framework designed for molecular optimization. ExLLM treats the LLM as the optimizer itself and introduces three complementary mechanisms: a single, evolving experience distilled from good and bad cases to avoid memory bloat; a $k$-offspring sampling scheme that leverages the autoregressive factorization to widen exploration per query; and a unified feedback adapter that integrates objectives, constraints and textual feedback for iterative prompting. The framework is simple to transfer with our task template custom evaluation functions and does not require any training.
@@ 3.1 Evolving Experience
Inspired by summary-style experience from ReEvo and ExpeL, we keep one continually updated experience $E_{t}$ into few hundreds of words. It captures actionable, transferable cues (e.g., frequently binding constraints, robust edit patterns, common invalid cases) without retrieving multiple entries. This avoids prompt bloat and reduces the risk of over-constraining exploration. From all historically evaluated candidates up to generation $t$, form an evidence set $D_{t}=G_{t}\cup B_{t}$, where $G_{t}$ are the top-$r$ candidates by fitness $F$ and $B_{t}$ are uniformly sampled from the lower-half of the fitness ranking. Let $S_{\theta}(\cdot)$ denote the same LLM used by the optimizer, prompted to produce a single concise memo (a few hundred words) that merges new evidence with prior experience. We update: $E_{t+1}\;=\;S_{\theta}\!\big(E_{t},\,D_{t})$ Operationally, $S_{\theta}$ folds in “good” and “bad” cases to reduce bias, promotes non-redundant micro-insights, and overwrites outdated content so that $E_{t}$ remains compact, general and up-to-date. At generation t+1t{+}1, $E_{t+1}$ is inserted directly into the LLM context without retrievals. To prevent over-exploitation, the experience is included in the prompt with probability $p_{\text{exp}}\in[0,1]$, sampled independently per call:
@@
This mechanism keeps prompts short on average, amortizes summarization cost (one update per generation), and maintains exploration headroom. We ablate $p_{\text{exp}}$ in Section 5 and observe that intermediate values yield the best trade-off between sample efficiency and diversity.
@@ 4 Experiment
Metrics
Objectives are normalized and direction-unified as in Sec. 3.3; fitness $F$ follows Eq. equation 1 with equal weights, consistent with Wang et al. (2024a). Hypervolume. Multi-objective coverage on the normalized (maximization) objective vectors using the reference point $\mathbf{r}=(1.1,\dots,1.1)$; larger values indicate better Pareto coverage. AUC Area under the curve of $F$ versus oracle evaluations up to budget $B$; this rewards methods that reach high values with fewer oracle calls (Gao et al., 2022). Validity: fraction passing RDKit parsing. Uniqueness: fraction of unique molecules proposed. Novelty: fraction absent from ZINC250K. Diversity: average pairwise Tanimoto distance between the Morgan fingerprints within the top-100. Efficiency LLM queries, API cost (USD), wall-clock time (hours) under the same budget.
Baselines
We compare against strong baselines spanning GA, BO, MCMC, RL, DL, and LLM-based methods: GB-GA, GB-BO, JT-VAE, MARS, REINVENT, MOLLEO, DyMol, and Genetic-GFN. For PMO-provided implementations (GB-GA, GB-BO, JT-VAE, MARS, REINVENT), we use the official PMO code and its default hyperparameters. For MOLLEO, DyMol, and Genetic-GFN, we use the authors’ public code with their default settings as documented in code/papers. For RL-based baselines (REINVENT, DyMol, Genetic-GFN), we use the scalar fitness $F$ as the reward signal during training.
@@ 4.1 Five-Objective Experiment Results
Top1 F 2.689 4.048 ± 0.114 3.838 ± 0.042 3.665 ± 0.129 3.891 ± 0.018 - 4.096 ± 0.155 - - 4.229 ± 0.050
Top10 F 2.683 4.019 ± 0.101 3.784 ± 0.027 3.647 ± 0.135 3.852 ± 0.019 - 4.044 ± 0.157 - - 4.186 ± 0.029
AUC-Top10 - 3.789 ± 0.079 3.712 ± 0.027 3.489 ± 0.104 3.740 ± 0.010 - 3.825 ± 0.102 - - 3.915 ± 0.010
Hypervolume 0.163 0.474 ± 0.190 0.364 ± 0.075 0.167 ± 0.050 0.488 ± 0.110 - 0.720 ± 0.172 - - 0.737 ± 0.038
Diversity 0.876 0.583 ± 0.032 0.847 ± 0.007 0.657 ± 0.033 0.826 ± 0.011 - 0.656 ± 0.111 - - 0.603 ± 0.055
Uniqueness - 0.786 ± 0.032 1.000 ± 0.000 1.000 ± 0.000 0.488 ± 0.128 - 0.672 ± 0.032 - - 0.829 ± 0.021
Validity - 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 - 0.930 ± 0.075 - - 0.937 ± 0.010
@@
Top1 F 4.329 4.583 ± 0.154 4.329 ± 0.000 4.605 ± 0.047 4.419 ± 0.074 - 4.699 ± 0.000 - - 4.699 ± 0.000
Top10 F 4.132 4.582 ± 0.167 4.132 ± 0.000 4.467 ± 0.066 4.181 ± 0.029 - 4.564 ± 0.064 - - 4.628 ± 0.043
AUC-Top10 - 4.130 ± 0.088 4.091 ± 0.000 4.237 ± 0.112 4.137 ± 0.028 - 4.362 ± 0.075 - - 4.481 ± 0.055
Hypervolume 0.917 0.968 ± 0.183 0.917 ± 0.000 1.275 ± 0.027 0.975 ± 0.041 - 1.168 ± 0.106 - - 1.175 ± 0.067
Diversity 0.793 0.424 ± 0.070 0.792 ± 0.001 0.630 ± 0.024 0.788 ± 0.002 - 0.600 ± 0.052 - - 0.491 ± 0.057
Uniqueness - 0.729 ± 0.041 1.000 ± 0.000 1.000 ± 0.000 0.432 ± 0.053 - 0.678 ± 0.000 - - 0.942 ± 0.006
Validity - 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 0.999 ± 0.000 - 0.913 ± 0.022 - - 0.790 ± 0.024
@@ 4.2 Results on PMO
Rank ($\downarrow$) 5 4 6 7 2 3 1
Table 3: Top-10 AUC of tasks in PMO (Gao et al., 2022) benchmark, including single-objective optimization and multi-objective optimization for 3 task types. ExLLM attains a total score of 19.165 (+7.3% improvement compared to the previous SOTA by MOLLEO), ranking first in 17/23 tasks and achieving the highest overall ranking.
@@ 4.3 Extended Experiments
We put the full tables and plots of the following experiments in the appendix. Cross-domain transfer. ExLLM achieves new records on circle packing and stellarator design, and shows strong performance on MOTSP/MOCVRP, SACS, NK2R peptide, and GCU operator design, with minimal task-specific changes. Constraint handling. We compare prompt-only penalties with promoting constraints to explicit objectives. The latter yields more stable optimization and better budgeted performance on tasks where constraint values are critical and variable.
5 Ablation Study
To fully characterize the framework, we ask: (1) how to determine $k$ and what’s the trade-off?; (2) how frequently should evolving experience be injected; (3) how does performance change with number of objectives; (4) how should constraints be handled for stable, budgeted gains? We hence quantify budgeted gains vs. local over-exploration by varying $k\!\in\!\{1,\dots,6\}$ in single-/multi-objective settings with two LLM backbones. And then identify when experience helps vs. hinders exploration by sweeping $p_{\text{exp}}\!\in\![0,1]$. We finally Tests scalability from 1 to 6 objectives including comparison with MOLLEO. Detailed tables and the $4^{th}$ ablation study are in the appendix.
Figure 2: $k$-offspring ablation studies. Top: single-objective optimization (JNK3). Bottom: five-objective optimization. The most consistent improvement on results and exploration occurs at kk=2 and 3, while higher $k$ may lead to local over-exploration.
$k$-offspring trade-offs Figure 2 shows the results of varying $k$ without using experience. For the five-objective plot, we subtract 3.5 from both AUC and $F$ before plotting to improve readability. Increasing $k$ improves exploratory breadth, especially in the single-objective JNK3 task, yet overly large $k$ tends to over-explore locally and limits global search under a fixed budget. Across tasks, kk=2 yields the largest and most stable gains in our sweep; kk=3 is less stable, and larger $k$ becomes increasingly unstable or degrades improvement.
$P_{exp}$ Top10 F AUC-Top10 Hypervolume Uniqueness Validity Diversity
0.0 4.245 ± 0.121 4.015 ± 0.074 0.850 ± 0.224 0.870 ± 0.014 0.900 ± 0.003 0.556 ± 0.106
0.3 4.260 ± 0.145 4.004 ± 0.077 0.881 ± 0.201 0.879 ± 0.014 0.905 ± 0.010 0.509 ± 0.057
0.5 4.301 ± 0.164 4.116 ± 0.040 0.905 ± 0.200 0.872 ± 0.015 0.908 ± 0.019 0.494 ± 0.032
0.7 3.986 ± 0.032 3.892 ± 0.007 0.555 ± 0.188 0.843 ± 0.016 0.930 ± 0.015 0.586 ± 0.070
0.9 4.030 ± 0.098 3.923 ± 0.051 0.571 ± 0.228 0.856 ± 0.020 0.926 ± 0.002 0.496 ± 0.021
Table 4: The ablation study of $P_{exp}$ shows that incorporating experience can notably improve performance and convergence, but it must be properly controlled to avoid restricting the exploration direction.
Experience Injection Our evolving experience is lightweight and general, yielding steady performance gains with only a modest impact on diversity. That said, in molecular optimization which has large discrete spaces, $P_{exp}$should be properly controlled to prevent over-conditioning the search and restricting exploration.
Figure 3: Experiment of different number of objectives.
[…]

### v5

@@ Abstract
Molecular design involves an enormous and irregular search space, where traditional optimizers such as Bayesian optimization, genetic algorithms, and generative models struggle to leverage expert knowledge or handle complex feedback. Recently, LLMs have been used as optimizers, achieving promising results on benchmarks such as PMO. However, existing approaches rely only on prompting or extra training, without mechanisms to handle complex feedback or maintain scalable memory. In particular, the common practice of appending or summarizing experiences at every query leads to redundancy, degraded exploration, and ultimately poor final outcomes under large-scale iterative search. We introduce ExLLM (Experience-Enhanced LLM optimization), an LLM-as-optimizer framework with three components: (1) a compact, evolving experience snippet tailored to large discrete spaces that distills non-redundant cues and improves convergence at low cost; (2) a simple yet effective k-offspring scheme that widens exploration per call and reduces orchestration cost; and (3) a lightweight feedback adapter that normalizes objectives for selection while formatting constraints and expert hints for iteration. ExLLM sets new state-of-the-art results on PMO and generalizes strongly—in our setup, it sets records on circle packing and stellarator design, and yields consistent gains across additional domains—requiring only a task-description template and evaluation functions to transfer. Code available at https://github.com/HuskyNian/ExLLM.
@@ 1 Introduction
We propose ExLLM, an LLM-as-optimizer framework designed for molecular optimization. ExLLM treats the LLM as the optimizer itself and introduces three complementary mechanisms: a single, evolving experience distilled from good and bad cases to avoid memory bloat; a $k$-offspring sampling scheme that leverages the autoregressive factorization to widen exploration per query; and a unified feedback adapter that integrates objectives, constraints and textual feedback for iterative prompting. The framework is simple to transfer: it works with only a task-description template and custom evaluation functions, and requires no training, as summarized in Figure 1.
@@ 2.1 Molecular Design with Machine Learning
Figure 1: ExLLM requires only two user inputs: a task-description template and evaluation functions, and operates without domain-specific training or tuning. The figure shows how a single unified workflow transfers seamlessly across different discrete problem optimization classes (chemical, geometric, combinatorial, physical, and code-generation tasks), achieving state-of-the-art or record-level performance under domain shift.
@@ 3.1 Evolving Experience
Prior work such as ReEvo or ExpeL uses summary-style memories for short-horizon reasoning. However, these designs are not suitable for large discrete optimization, where redundancy accumulation and exploration collapse are major problems. Here, we keep one continually updated experience $E_{t}$ into few hundreds of words. It captures actionable, transferable cues (e.g., frequently binding constraints, robust edit patterns, common invalid cases) without retrieving multiple entries. This avoids prompt bloat and reduces the risk of over-constraining exploration. From all historically evaluated candidates up to generation $t$, form an evidence set $D_{t}=G_{t}\cup B_{t}$, where $G_{t}$ are the top-$r$ candidates by fitness $F$ and $B_{t}$ are uniformly sampled from the lower-half of the fitness ranking. Let $S_{\theta}(\cdot)$ denote the same LLM used by the optimizer, prompted to produce a single concise memo (a few hundred words) that merges new evidence with prior experience. We update: $E_{t+1}\;=\;S_{\theta}\!\big(E_{t},\,D_{t})$ Operationally, $S_{\theta}$ folds in “good” and “bad” cases to reduce bias, promotes non-redundant micro-insights, and overwrites outdated content so that $E_{t}$ remains compact, general and up-to-date. At generation t+1t{+}1, $E_{t+1}$ is inserted directly into the LLM context without retrievals. To prevent over-exploitation, the experience is included in the prompt with probability $p_{\text{exp}}\in[0,1]$, sampled independently per call:
@@
This mechanism keeps prompts short on average, amortizes summarization cost (one update per generation), and maintains exploration headroom. We ablate $p_{\text{exp}}$ in Section 6.2 and observe that intermediate values yield the best trade-off between sample efficiency and diversity.
@@ 4 Experiment
Metrics Objectives are normalized and direction-unified as in Sec. 3.3; fitness $F$ follows Eq. equation 1 with equal weights, consistent with Wang et al. (2024a). Hypervolume. Multi-objective coverage on the normalized (maximization) objective vectors using the reference point $\mathbf{r}=(1.1,\dots,1.1)$; larger values indicate better Pareto coverage. AUC Area under the curve of $F$ versus oracle evaluations up to budget $B$; this rewards methods that reach high values with fewer oracle calls (Gao et al., 2022). Validity: fraction passing RDKit parsing. Uniqueness: fraction of unique molecules proposed. Novelty: fraction absent from ZINC250K. Diversity: average pairwise Tanimoto distance between the Morgan fingerprints within the top-100. Efficiency LLM queries, API cost (USD), wall-clock time (hours) under the same budget.
Baselines We compare against strong baselines spanning GA, BO, MCMC, RL, DL, and LLM-based methods: GB-GA, GB-BO, JT-VAE, MARS, REINVENT, MOLLEO, DyMol, and Genetic-GFN. For PMO-provided implementations (GB-GA, GB-BO, JT-VAE, MARS, REINVENT), we use the official PMO code and its default hyperparameters. For MOLLEO, DyMol, and Genetic-GFN, we use the authors’ public code with their default settings as documented in code/papers. For RL-based baselines (REINVENT, DyMol, Genetic-GFN), we use the scalar fitness $F$ as the reward signal during training.
@@ 4.1 Five-Objective Experiment Results
Top1 F 2.689 4.048 ± 0.114 3.838 ± 0.042 3.665 ± 0.129 3.891 ± 0.018 N/A 4.096 ± 0.155 N/A N/A 4.229 ± 0.050
Top10 F 2.683 4.019 ± 0.101 3.784 ± 0.027 3.647 ± 0.135 3.852 ± 0.019 N/A 4.044 ± 0.157 N/A N/A 4.186 ± 0.029
AUC-Top10 N/A 3.789 ± 0.079 3.712 ± 0.027 3.489 ± 0.104 3.740 ± 0.010 N/A 3.825 ± 0.102 N/A N/A 3.915 ± 0.010
Hypervolume 0.163 0.474 ± 0.190 0.364 ± 0.075 0.167 ± 0.050 0.488 ± 0.110 N/A 0.720 ± 0.172 N/A N/A 0.737 ± 0.038
Diversity 0.876 0.583 ± 0.032 0.847 ± 0.007 0.657 ± 0.033 0.826 ± 0.011 N/A 0.656 ± 0.111 N/A N/A 0.603 ± 0.055
Uniqueness N/A 0.786 ± 0.032 1.000 ± 0.000 1.000 ± 0.000 0.488 ± 0.128 N/A 0.672 ± 0.032 N/A N/A 0.829 ± 0.021
Validity N/A 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 N/A 0.930 ± 0.075 N/A N/A 0.937 ± 0.010
@@
Top1 F 4.329 4.583 ± 0.154 4.329 ± 0.000 4.605 ± 0.047 4.419 ± 0.074 N/A 4.699 ± 0.000 N/A N/A 4.699 ± 0.000
Top10 F 4.132 4.582 ± 0.167 4.132 ± 0.000 4.467 ± 0.066 4.181 ± 0.029 N/A 4.564 ± 0.064 N/A N/A 4.628 ± 0.043
AUC-Top10 N/A 4.130 ± 0.088 4.091 ± 0.000 4.237 ± 0.112 4.137 ± 0.028 N/A 4.362 ± 0.075 N/A N/A 4.481 ± 0.055
Hypervolume 0.917 0.968 ± 0.183 0.917 ± 0.000 1.275 ± 0.027 0.975 ± 0.041 N/A 1.168 ± 0.106 N/A N/A 1.175 ± 0.067
Diversity 0.793 0.424 ± 0.070 0.792 ± 0.001 0.630 ± 0.024 0.788 ± 0.002 N/A 0.600 ± 0.052 N/A N/A 0.491 ± 0.057
Uniqueness N/A 0.729 ± 0.041 1.000 ± 0.000 1.000 ± 0.000 0.432 ± 0.053 N/A 0.678 ± 0.000 N/A N/A 0.942 ± 0.006
Validity N/A 1.000 ± 0.000 1.000 ± 0.000 1.000 ± 0.000 0.999 ± 0.000 N/A 0.913 ± 0.022 N/A N/A 0.790 ± 0.024
@@ 4.2 Results on PMO
Rank ($\downarrow$) 7 5 8 9 2 4 1
Table 3: Top-10 AUC of tasks in PMO (Gao et al., 2022) benchmark, including single-objective optimization and multi-objective optimization for 3 task types. ExLLM attains a total score of 19.165 (+7.3% improvement compared to the previous SOTA by MOLLEO), ranking first in 17/23 tasks and achieving the highest overall ranking. The models with 3rd and 6th totao score are in the additional LLM-based results in Appendix 6.3.
@@ 4.3 Extended Experiments
Cross-domain transfer. To fully demonstrate the generalizability of ExLLM, we conduct experiements in multiple domains, including circle packing (mathematics), stellarator coil optimization (physics), and offshore jacket structural analysis (offshore structural engineering), MOTSP and MOCVRP (classical optimization problems), NK2R peptide design (peptide-drug discovery), GCU operator optimization (high performance computing kernel design). ExLLM achieves new records on circle packing and stellarator design benchmark (ConStellaration Cadena et al. (2025)), and achieves SOTA performance on MOCVRP and offshore jacket structural optimization, and competitive performance on MOTSP, peptide deisgn and kernel optimziation.
Circle packing in a unit square Given an integer $n$, the task is to place $n$ congruent circles inside the unit square while maximizing the common radius $r$ subject to two strict feasibility constraints: (i) no overlaps, $\lVert c_{i}-c_{j}\rVert_{2}\geq 2r$, and (ii) all circle centers must satisfy $r\leq x_{i},y_{i}\leq 1-r$. The optimization landscape is very unstable, sub-millimeter changes in coordinates easily break feasibility, and the input state $(x_{i},y_{i},r)$ provides no structural cues for the model. Only a single scalar objective (the sum of radius) is returned, making gradient-free global search difficult. As shown in Table 8, ExLLM attains new best-known radii for $n=26$ and $n=32$ and matches all published records for n=27n=27–31. Existing records for these middle cases are available only to three decimal places, so it is unclear whether ExLLM further improves beyond that precision. These results are achieved with the same hyperparameters used in all other domains, demonstrating ExLLM’s ability to solve highly sensitive geometric–continuous problems under strict feasibility constraints.
Table 4: MOCPOP benchmark. ExLLM achieves SOTA on MOCVRP and ranks highly on MOTSP.
Method MOTSP↑ MOCVRP↑
Pymoo 0.983 0.956
ReEvo 1.029 1.035
AIDE 1.021 1.006
FunSearch 1.023 1.032
AlphaEvolve 1.0293 1.0318
ExLLM 1.0273 1.0704
Table 5: Offshore platform jackets optimization. ExLLM finds the lightest feasible design.
Method Weight↓ Stress↓
Human 218.0 0.024
GA 32.24 0.093
MOEAD 37.76 0.137
RS 49.33 0.435
Ablation 14.90 0.814
ExLLM 13.60 0.508
Table 6: Stellarator optimization (ConStellaration). ExLLM exceeds official SOTA on P2/P3.
Method P2↑ P3↑
trust-constr Fail Fail
COBYQA Fail Fail
ALM–NGOpt 0.431 129.796 (97)
ExLLM 0.505 133.634 (103)
Table 7: Circle packing (n=26–32), rotated for compact layout. ExLLM improves all best-known results.
n Record ExLLM
26 2.635977 (AlphaEvolve (Novikov et al., 2025a)) 2.635983
27 2.685+ (Friedman, 2025) 2.68598
28 2.737+ (Friedman, 2025) 2.73740
29 2.790+ (Friedman, 2025) 2.79034
30 2.842+ (Friedman, 2025) 2.84267
31 2.889+ (Friedman, 2025) 2.88997
32 2.937+ (AlphaEvolve (Novikov et al., 2025a)) 2.939+
Table 8: NK2R peptide design (AlphaFold3 ipTM). ExLLM peptides outperform natural ligand NKA.
Peptide ipTM(NKA) ipTM(ours) $\Delta$
P1 0.09 0.87 0.78
P2 0.10 0.88 0.78
P3 0.12 0.88 0.76
P4 0.12 0.88 0.76
P5 0.11 0.86 0.75
P6 0.10 0.84 0.74
P7 0.13 0.87 0.74
P8 0.14 0.88 0.74
P9 0.11 0.85 0.74
P10 0.13 0.87 0.74
Stellarator Design The ConStellaration benchmark defines two quasi-isodynamic (QI) stellarator optimization tasks: (P2) maximizing a coil-friendly QI score under five physics constraints, and (P3) jointly optimizing MHD stability and coil simplicity under the five physics constraints as well. A design is scored only if all constraints are satisfied; otherwise, its score is zero. The feasible region is extremely narrow, and small geometric deviations can simultaneously violate multiple plasma-physics constraints. Standard baselines (trust-constr, COBYQA) fail to find any feasible point. As shown in Table 8, ExLLM consistently identifies feasible solutions and achieves new SOTA on both tasks: a 17% improvement on P2 (0.505 vs. 0.431), a 3% improvement on P3 (133.634 vs. 129.796) and a 6% improvement on the single point(97 vs. 103) on P3. This demonstrates that ExLLM can navigate fragile multi-constraint physics landscapes without any domain-specific tuning.
Offshore jacket optimization The SACS benchmark requires optimizing offshore jacket structures under three competing engineering goals: minimizing total weight, while satisfying axial-load and bending/torsional stress limits. Each candidate is evaluated by commercial SACS finite-element analysis, which provides only black-box outputs (weight and stress ratios). The search space mixes discrete structural-section choices with continuous dimensions, and SACS evaluations are costly, making sample efficiency critical. Table 8 shows that ExLLM finds a feasible design weighing only 13.6 tons, more than 93% lighter than the 218-ton human baseline, while satisfying all stress constraints. It also outperforms GA, MOEAD, and random search. This highlights ExLLM’s capacity to handle mixed-variable engineering optimization with expensive, black-box oracles.
MOCPOP The Multi-Objective Traveling Salesman Problem (MOTSP) and the Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP) tasks involve large-scale ($n=100$) routing under strict permutation and capacity constraints. Both problems exhibit strong multi-objective conflict: small changes in coordinates or routes can drastically alter feasibility or the relative objective ordering, and scalarized heuristics often collapse the Pareto front. As reported in Table 8, ExLLM achieves new SOTA hypervolume on MOCVRP and ranks second only to AlphaEvolve on MOTSP, outperforming strong baselines such as Pymoo, ReEvo, AIDE, and FunSearch. These results show that ExLLM effectively reasons over discrete combinatorial structures and generates high-quality, feasible offspring without relying on brittle crossover heuristics.
NK2R peptide design The NK2R task requires designing peptides that (i) share $30\%$ sequence similarity with the native ligand NKA and (ii) achieve higher AlphaFold3 complex ipTM with NK2R. Because similarity rules, formatting constraints, and ipTM comparison to NKA jointly determine success, most candidates score zero and the landscape is extremely sparse. Table 8 shows that ExLLM produces many peptides surpassing NKA within 1,000 evaluations, with best margins around +0.78. All reported peptides satisfy the similarity and length constraints. This indicates that ExLLM can perform constrained molecular design using only AF3 black-box feedback.
GCU operator design This domain evaluates whether ExLLM can generate high-performance GPU kernels for Tencent’s GCU across three operators (Var, SiLU, GEMM v2). The task is challenging due to a new programming model, scarce public examples, strict compiler requirements, and 5–10 minute compile–run cycles where naive LLM code fails $90\%$ of the time. With only 300 evaluations per operator, ExLLM raises compile success from $10\%$ to about 85%, produces fully correct kernels, and significantly improves execution latency. These submissions placed in the top-10 of the Tencent Kaiwu 2025 competition qualifier and earned a second prize award, with all submitted kernels fully LLM-generated and no manual code edits. This demonstrates ExLLM’s ability to operate in real-world domain-shifted, performance-critical code-generation settings.